OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 6 4/5 inches and a width of 2 2/3 inches?

OpenStudy (lochana):

6 4/5 = 34/5 2 2/3 = 8/3 

OpenStudy (lochana):

perimeter = 34/5 + 34/5 + 8/3 + 8/3 =71/15

OpenStudy (lochana):

and final answer = 4 11/15 inches

OpenStudy (anonymous):

thanks so much!

OpenStudy (lochana):

you are welcome!